# The purpose of an estimator is what it does: Misspecification, estimands, and over-identification

### v4

@@ 3. Choice of Estimator and Estimand
Consider estimators $\hat{\theta}(X^{n})\in\Theta$ that are consistent under the model, i.e., for all $(\theta,P)\in\mathcal{M},$ $\hat{\theta}\overset{p}{\longrightarrow}\theta$ as $n\to\infty$. Under standard regularity conditions on the nesting model $\mathcal{M}^{*}$, $\hat{\theta}$ still converges in probability to some limit when $P\in\mathcal{P}^{*}\setminus\mathcal{P}$. Denote this limit by $\text{plim}(\hat{\theta},P).$ Because $\hat{\theta}$ is consistent under the model, $\text{plim}(\hat{\theta},P)=\theta^{*}(P)$ for every $P\in\mathcal{P}$. However, different choices of estimator will typically imply different limits $\text{plim}(\hat{\theta},P)$ for $P$ outside the initial model, $P\in\mathcal{P}^{*}\setminus\mathcal{P}$. [^12] Indeed, as noted by Chen and Santos (2018) there is a close connection between testability of a statistical model $\mathcal{P}$ and the existence of distinct estimators for statistical functionals defined on the model. Thus, when choosing which estimator to use under the model $\mathcal{M},$ we are also (implicitly) choosing which parameter we wish to estimate when the model is statistically misspecified.
@@
Under the model $\mathcal{M}$, we may estimate $\theta$ using the ordinary least squares (OLS) estimator $\hat{\theta}_{\text{OLS}}=\frac{\operatorname{Cov}_{n}(W_{i},Y_{i})}{\operatorname{Var}_{n}(W_{i})},$ for $\operatorname{Cov}_{n}$ and $\operatorname{Var}_{n}$ the sample covariance and variance, respectively. Alternatively, we may instead estimate $\theta$ using the weighted least squares (WLS) estimator $\hat{\theta}_{\text{WLS}}=\frac{\operatorname{Cov}_{n,\omega}(W_{i},Y_{i})}{\operatorname{Var}_{n,\omega}(W_{i})},$ where $\operatorname{Cov}_{n,\omega}$ and $\operatorname{Var}_{n,\omega}$ weight observation $i$ by $\omega_{i}=\omega(W_{i})$ for a weight function $\omega.$ By the Gauss–Markov theorem, the efficient estimator uses $\omega_{i}=\operatorname{Var}_{P}(Y_{i}\mid W_{i})^{-1}$. Consistency implies that $\text{plim}(\hat{\theta}_{\text{OLS}},P)=\text{plim}(\hat{\theta}_{\text{WLS}},P)$ for all $P\in\mathcal{P}.$ By contrast, under statistical misspecification $\text{plim}(\hat{\theta}_{\text{OLS}},P)=\frac{\operatorname{Cov}_{P}(W_{i},Y_{i})}{\operatorname{Var}_{P}(W_{i})}$ while $\text{plim}(\hat{\theta}_{\text{WLS}},P)=\frac{\operatorname{Cov}_{P,\omega}(W_{i},Y_{i})}{\operatorname{Var}_{P,\omega}(W_{i})},$ where these limits can differ for $P\not\in\mathcal{P}.$ Hence, once we allow for misspecification WLS is not more efficient than OLS: the two simply estimate different things. $\blacksquare$
@@
For instance, the two-step efficient GMM estimator, which minimizes the large-sample variance of the estimate under the model, sets $\hat{\Omega}$ equal to the inverse of an estimate of the asymptotic variance of $\sqrt{n}g_{n}(\psi).$
If we define $\Omega_{P}=\text{plim}(\hat{\Omega},P)$—assumed to exist—and $g_{P}(\psi)=\mathbb{E}_{P}[g(X_{i},\psi)],$ then under mild conditions (see Imbens (1997); Hall and Inoue (2003)),
@@
When the model is statistically correct the argmin in the GMM objective does not depend on the weighting matrix, and thus neither does the limit of $\hat{\theta}_{\Omega}$. In contrast, when the model is statistically misspecified different weighting matrices yield different estimands. Hence, under misspecification the two-step GMM estimator is no more efficient than any other estimator: Once again each weighting leads us to recover a different parameter. $\blacksquare$
@@
Previously, the estimators we discussed differed only by how they weighted a common set of objects ($W_{i}$ values in one case and moments in the other). In other settings, estimators exploit the data in qualitatively different ways. To illustrate, suppose we have a parametric model which implies that the data are drawn from $P_{\psi}$, for a finite-dimensional parameter $\psi,$ where $\theta=\vartheta(\psi)$. Assume $P_{\psi}$ is a continuous distribution with density $p_{\psi}$. Our econometric model is thus $\mathcal{M}=\left\{(\vartheta(\psi),P_{\psi}):\psi\in\Psi\right\},$ while our statistical model is $\mathcal{P}=\left\{P_{\psi}:\psi\in\Psi\right\}.$
Under the model $\mathcal{M}$ we may estimate $\theta$ by maximum likelihood, taking $\hat{\psi}_{\text{MLE}}=\operatorname*{arg\,max}_{\psi\in\Psi}\log p_{\psi}(X^{n})$ and $\hat{\theta}_{\text{MLE}}=\vartheta(\hat{\psi}_{\text{MLE}}).$ Alternatively, we may estimate the model by the method of simulated moments, choosing a set of $\mathbb{R}^{k}$-valued functions $h$, constructing moments comparing the realized values of $h(X_{i})$ to their model-implied means $g(X_{i},\psi)=h(X_{i})-\mathbb{E}_{P_{\psi}}[h(X_{i})],$ and computing GMM estimates $\hat{\psi}_{\Omega}$ and $\hat{\theta}_{\Omega}$ for some choice of weighting matrix. Provided the moment equations have a unique solution under $P_{\psi}$ for all $\psi\in\Psi,$ these estimators have the same probability limit under the model but maximum likelihood will be (weakly) more efficient.
As previously discussed, under misspecification and provided regularity conditions hold, the GMM estimator $\hat{\psi}_{\Omega}$ converges to the value which minimizes the $\Omega_{P}$-weighted deviation of the moment conditions from zero. In contrast, $\text{plim}(\hat{\psi}_{\text{MLE}},P)=\operatorname*{arg\,max}_{\psi\in\Psi}\mathbb{E}_{P}[\log p_{\psi}(X^{n})],$ which minimizes the Kullback–Leibler divergence between the distribution of the data and the model-implied distribution (White, 1982). $\blacksquare$
@@
The results of Yitzhaki (1996) imply that the least-squares estimand $\text{plim}(\hat{\theta}_{OLS},P)$ can be written as a weighted average of the partial derivative of $\mathbb{E}_{P}[Y_{i}\mid W_{i}]$ with respect to $W_{i},$
$\text{plim}(\hat{\theta}_{OLS},P)=\mathbb{E}_{P}\left[\eta_{P}(W_{i})\frac{\partial}{\partial w}\mathbb{E}_{P}[Y_{i}\mid W_{i}]\right]$
for particular weights $\eta_{P}(W_{i})\geq 0$ with $\mathbb{E}_{P}[\eta_{P}(W_{i})]=1.$ If the nesting model $\mathcal{M}^{*}$ maintains random assignment and differentiability of potential outcomes $Y_{i}(w),$ it follows that $\text{plim}(\hat{\theta}_{OLS},P)=\mathbb{E}_{P}[\eta_{P}(W_{i})\frac{\partial}{\partial w}Y_{i}(W_{i})],$ so OLS recovers a weighted average of causal effects, though typically not the unweighted average specified by $\theta.$ One can show an analogous result for weighted least squares, so the choice between the OLS and WLS estimands reduces to preference between the two sets of weights. $\blacksquare$
@@
In some economic applications of parametric models, researchers appear to prefer the estimates from simulated method of moments to maximum likelihood. [^14] For instance, Bordalo et al. (2020) write “We prefer this method to maximum likelihood for two reasons. First, one advantage of our model is that it is simple and transparent. However, this simplicity comes at the cost of likely misspecification, and it is well known that with misspecification concerns moment estimators are often more reliable.” This view has a long history in economics, with many researchers preferring SMM or related “calibration” procedures to maximum likelihood because of the ability to select moments that, in their judgment, more directly reflect the economic quantity of interest—see Dawkins et al. (2001) for an overview of calibration and extensive references.
It is natural to interpret these statements as (implicit) restrictions on the nesting model researchers have in mind, which ensure that some estimands are more closely related to the parameters of interest that others. Such restrictions are rarely stated formally, however. Since these considerations appear to actively guide estimation choices, more explicit articulation of the rationale for a given estimator is useful, and formalizing such arguments seems a potentially fruitful area for future econometric research. $\blacksquare$
@@
As with OLS, the literature has characterized the GMM estimand for specific choices of moments, weights, and nesting models. For example, Imbens and Angrist (1994) and Angrist and Imbens (1995) show that, under monotonicity, two-stage least squares—a special case of GMM—recovers a positive-weight average of treatment effects. In finance, Hansen and Jagannathan (1997) show that in a misspecified asset pricing model the GMM objective with a specific weighting matrix measures the squared error in approximating an admissible stochastic discount factor, and so measures the “size” of pricing errors in a precise sense. $\blacksquare$
@@
As our discussion of parametric models suggests, even beyond settings covered by existing theoretical guarantees, researchers’ choice of weighting matrices for GMM and related methods (e.g. simulated method of moments, minimum distance) are often informed by concerns about misspecification. To systematically explore this, we reviewed empirical papers published in the American Economic Review between 2020 and 2024 which a keyword search suggested used moment-based methods (excluding methods like ordinary least squares, two-stage least squares, and maximum likelihood which may be cast as special cases of GMM but which the authors did not describe as such). The search turned up 36 papers. We refer to these papers collectively as our “GMM sample.” In this sample, 22 of the papers were over-identified, and in 14 of these no results were reported based on an efficient weighting matrix. [^15] In addition, some of the papers which notionally use efficient weights plug in parameters estimated in a first step. This changes the form of the “correct” optimal weighting matrix, but is not accounted for in the implemented weighting matrices. Consequently, the number of papers in fact using theoretically efficient weights is still lower. For those models where results based on efficient GMM are not reported, the most common (non-exclusive) choices were (i) the inverse of the diagonal of the asymptotic variance-covariance matrices (4 papers) (ii) identity weighting matrix (6 papers), and (iii) user-selected weighting matrices, often explicitly chosen based on domain knowledge (7 papers).
@@
One reason the “optimal” weighting matrix may not be used in some cases is that in settings where the sample size is small relative to the number of moments, it is known that estimation of the optimal weighting matrix can lead to bias in both the estimate and the standard errors (see e.g. Altonji and Segal, 1996; Windmeijer, 2005). One suggestion in such cases is to use the inverse of the diagonal alone, which Altonji and Segal (1996) found yielded “a substantial bias reduction” their problem. [^16] Consistent with this intepretation, Abebe et al. (2021a) write that “In line with the recent literature, we use this simple weighting matrix instead of the theoretically optimal weighting matrix, which may suffer from small sample bias (Altonji and Segal, 1996).” However, together with optimal weighting this still covers only about half of the over-identified papers in our GMM sample.
@@
For the remaining two choices, identity weights and hand-selected weights, researchers are directly choosing the weight assigned to different moments, and thus the estimand. This is most obvious for hand-selected weighting matrices, where in many cases the authors explicitly motivate their choice of weighting matrix based on their goals for estimation. [^17] For instance, Antràs and De Gortari (2020) write that “In order to guarantee that our model provides a proper quantitative evaluation of the general-equilibrium workings of the world economy, we place a higher weight on matching the empirical moments of larger trade flows.” Similarly, Einav et al. (2020) write that “Because identification is primarily driven by the discontinuous change in screening rates at age 40, we weight more heavily moments that are closer to age 40 than moments that are associated with younger and older ages.” Perhaps more subtly, for identity weighting matrices the scale that the authors choose for each moment (e.g. whether to measure a monetary outcome in dollars or thousands of dollars) determines the weight. [^18] Weighting by the inverse of the diagonal of the variance eliminates dependence on the scale of individual moments, but is still sensitive to more general linear transformations of the moments, for instance replacing control and treatment means by control means and treatment-control contrasts. In Lise and Postel-Vinay (2020), for instance, the authors present a standard error formula based on identity weighting (see their Footnote 31), while their code uses a hand-selected diagonal but non-identity weighting (which is consistent with the discussion in their footnote under an appropriate choice of units for the moments).
@@
A growing literature characterizes estimands obtained from misspecified models in additional settings. An early positive result by Angrist (1998) shows that in linear models where a binary treatment is as good as randomly assigned given a discrete covariate, regressions on the treatment and saturated dummies for the covariate recover a positive weighted average of causal effects. More recently, however, a large literature warns that estimands from linear (regression or instrumental variables) models fail to recover positively weighted averages of causal effects in important cases. Examples include difference in difference and event-study models estimated via two-way fixed effect regressions (de Chaisemartin and D’Haultfœuille, 2020; Callaway and Sant’Anna, 2021; Sun and Abraham, 2021), IV models estimated by two stage least squares regressions with multiple endogenous variables or insufficiently rich controls (Słoczyński, 2020; Bhuller and Sigstad, 2024; Blandhol et al., 2025), nonlinear estimators for linear IV, such as LIML and continuously updating GMM (Kolesar, 2013; Andrews, 2019), and linear regressions including both controls and multiple treatments (Goldsmith-Pinkham et al., 2024). Chen (2025) develops a general method for characterizing the causal estimand, if any, recovered by linear regression estimates.
Other recent papers characterize estimands in nonlinear models, including Andrews et al. (2025) who study a class of nonlinear structural models and provide necessary and sufficient conditions for instrumental variables estimates to recover linear transformations of causal effects, and Kolesár and Plagborg-Møller (2025) who consider estimation of causal effects in time-series models and derive positive results for certain linear estimators, along with negative results for some popular nonlinear estimators. Other recent work extends the concept of a pseudo-true value to set-identified models (Andrews and Kwon, 2023; Kaido and Molinari, 2024). $\blacksquare$

### v5

@@ 3. Choice of Estimator and Estimand
Consider estimators $\hat{\theta}(X^{n})\in\Theta$ that are consistent under the model, i.e., for all $(\theta,P)\in\mathcal{M},$ $\hat{\theta}\overset{p}{\longrightarrow}\theta$ as $n\to\infty$. Under standard regularity conditions on the nesting model $\mathcal{M}^{*}$, $\hat{\theta}$ still converges in probability to some limit when $P\in\mathcal{P}^{*}\setminus\mathcal{P}$. Denote this limit by $\text{plim}(\hat{\theta},P).$ Because $\hat{\theta}$ is consistent under the model, $\text{plim}(\hat{\theta},P)=\theta^{*}(P)$ for every $P\in\mathcal{P}$. However, different choices of estimator will typically imply different limits $\text{plim}(\hat{\theta},P)$ for $P$ outside the initial model, $P\in\mathcal{P}^{*}\setminus\mathcal{P}$. [^13] Indeed, as noted by chen2018overidentification there is a close connection between testability of a statistical model $\mathcal{P}$ and the existence of distinct estimators for statistical functionals defined on the model. Thus, when choosing which estimator to use under the model $\mathcal{M},$ we are also (implicitly) choosing which parameter we wish to estimate when the model is statistically misspecified.
@@
Under the model $\mathcal{M}$, we may estimate $\theta$ using the ordinary least squares (OLS) estimator $\hat{\theta}_{\text{OLS}}=\frac{\operatorname{Cov}_{n}(W_{i},Y_{i})}{\operatorname{Var}_{n}(W_{i})},$ for $\operatorname{Cov}_{n}$ and $\operatorname{Var}_{n}$ the sample covariance and variance, respectively. Alternatively, we may instead estimate $\theta$ using the weighted least squares (WLS) estimator $\hat{\theta}_{\text{WLS}}=\frac{\operatorname{Cov}_{n,\omega}(W_{i},Y_{i})}{\operatorname{Var}_{n,\omega}(W_{i})},$ where $\operatorname{Cov}_{n,\omega}$ and $\operatorname{Var}_{n,\omega}$ weight observation $i$ by $\omega_{i}=\omega(W_{i})$ for a weight function $\omega.$ By the Gauss–Markov theorem, the efficient estimator uses $\omega_{i}=\operatorname{Var}_{P}(Y_{i}\mid W_{i})^{-1}$. Consistency implies that $\text{plim}(\hat{\theta}_{\text{OLS}},P)=\text{plim}(\hat{\theta}_{\text{WLS}},P)$ for all $P\in\mathcal{P}.$ By contrast, under statistical misspecification $\text{plim}(\hat{\theta}_{\text{OLS}},P)=\frac{\operatorname{Cov}_{P}(W_{i},Y_{i})}{\operatorname{Var}_{P}(W_{i})}$ while $\text{plim}(\hat{\theta}_{\text{WLS}},P)=\frac{\operatorname{Cov}_{P,\omega}(W_{i},Y_{i})}{\operatorname{Var}_{P,\omega}(W_{i})},$ where these limits can differ for $P\not\in\mathcal{P}.$ Hence, once we allow for misspecification, WLS is not more efficient than OLS: The two simply estimate different things. $\blacksquare$
@@
For instance, the two-step efficient GMM estimator, which minimizes the large-sample variance of the estimate under the model, sets $\hat{\Omega}$ equal to the inverse of an estimate of the asymptotic variance of $\sqrt{n}g_{n}(X^{n},\psi).$
If we define $\Omega_{P}=\text{plim}(\hat{\Omega},P)$—assumed to exist—and $g_{P}(\psi)=E_{P}[g(X_{i},\psi)],$ then under mild conditions (see imbens1997one; hall2003large),
@@
When the model is statistically correct, the argmin of the population GMM objective does not depend on the weighting matrix, and thus neither does the limit of $\hat{\theta}_{\Omega}$. In contrast, when the model is statistically misspecified different weighting matrices yield different estimands. Hence, under misspecification the two-step GMM estimator is no more efficient than any other estimator: Once again each weighting leads us to recover a different parameter. $\blacksquare$
@@
Previously, the estimators we discussed differed only by how they weighted a common set of objects ($W_{i}$ values in one case and moments in the other). In other settings, estimators exploit the data in qualitatively different ways. To illustrate, suppose we have a parametric model which implies that the data are drawn from $P_{\psi}$, for a finite-dimensional parameter $\psi,$ where $\theta=\vartheta(\psi)$. Assume $P_{\psi}$ is a continuous distribution with density $p_{\psi}$, and that $X_{i}$ is identically distributed. Our econometric model is thus $\mathcal{M}=\left\{(\vartheta(\psi),P_{\psi}):\psi\in\Psi\right\},$ while our statistical model is $\mathcal{P}=\left\{P_{\psi}:\psi\in\Psi\right\}.$
Under the model $\mathcal{M}$ we may estimate $\theta$ by maximum likelihood, taking
$\hat{\psi}_{\text{MLE}}=\operatorname*{arg\,max}_{\psi\in\Psi}\log p_{\psi}(X^{n})$
and $\hat{\theta}_{\text{MLE}}=\vartheta\left(\hat{\psi}_{\text{MLE}}\right).$ Alternatively, we may estimate the model by the method of simulated moments, choosing a set of $\mathbb{R}^{k}$-valued functions $h$, constructing moments comparing the realized values of $h(X_{i})$ to their model-implied means, $g(X_{i},\psi)=h(X_{i})-E_{P_{\psi}}[h(X_{i})],$ and computing GMM estimates $\hat{\psi}_{\Omega}$ and $\hat{\theta}_{\Omega}$ for some choice of weighting matrix. Provided the moment equations have a unique solution under $P_{\psi}$ for all $\psi\in\Psi,$ these estimators have the same probability limit under the model but maximum likelihood will be (weakly) more efficient.
As previously discussed, under misspecification and provided regularity conditions hold, the GMM estimator $\hat{\psi}_{\Omega}$ converges to the value which minimizes the $\Omega_{P}$-weighted deviation of the moment conditions from zero. In contrast, $\text{plim}(\hat{\psi}_{\text{MLE}},P)=\lim_{n\to\infty}\operatorname*{arg\,max}_{\psi\in\Psi}E_{P}[\log p_{\psi}(X^{n})],$ which minimizes the Kullback–Leibler divergence between the distribution of the data and the model-implied distribution (white1982maximum). $\blacksquare$
@@
The results of yitzhaki1996using imply that the least-squares estimand $\text{plim}(\hat{\theta}_{\mathrm{OLS}},P)$ can be written as a weighted average of the partial derivative of $E_{P}[Y_{i}\mid W_{i}]$ with respect to $W_{i},$
$\text{plim}(\hat{\theta}_{\mathrm{OLS}},P)=E_{P}\left[\eta_{P}(W_{i})\frac{\partial}{\partial w}E_{P}[Y_{i}\mid W_{i}]\right]$
for particular weights $\eta_{P}(W_{i})\geq 0$ with $E_{P}[\eta_{P}(W_{i})]=1.$ If the nesting model $\mathcal{M}^{*}$ maintains random assignment and differentiability of potential outcomes $Y_{i}(w),$ it follows that $\text{plim}(\hat{\theta}_{\mathrm{OLS}},P)=E_{P}[\eta_{P}(W_{i})\frac{\partial}{\partial w}Y_{i}(W_{i})],$ so OLS recovers a weighted average of causal effects, though typically not the unweighted average specified by $\theta.$ One can show an analogous result for weighted least squares, so the choice between the OLS and WLS estimands reduces to preference between the two sets of weights. $\blacksquare$
@@
In some economic applications of parametric models, researchers appear to prefer the estimates from simulated method of moments to maximum likelihood. [^15] For instance, bordalo2020overreaction write “We prefer this method to maximum likelihood for two reasons. First, one advantage of our model is that it is simple and transparent. However, this simplicity comes at the cost of likely misspecification, and it is well known that with misspecification concerns moment estimators are often more reliable.” This view has a long history in economics, with many researchers preferring SMM or related “calibration” procedures to maximum likelihood because of the ability to select moments that, in their judgment, more directly reflect the economic quantity of interest—see dawkins2001calibration for an overview of calibration and extensive references.
It is natural to interpret these statements as (implicit) restrictions on the nesting model researchers have in mind, which ensure that some estimands are more closely related to the parameters of interest than others or, relatedly, that some estimation criteria better reflect the researcher’s objectives. Such restrictions are rarely stated formally, however. Since these considerations appear to actively guide estimation choices, more explicit articulation of the rationale for a given estimator is useful, and formalizing such arguments seems a potentially fruitful area for future econometric research. $\blacksquare$
@@
As with OLS, the literature has characterized the GMM estimand for specific choices of moments, weights, and nesting models. For example, 209996d6-9cfc-38cf-a7bb-28a4c5d811aa and angrist1995two show that, under monotonicity, two-stage least squares—a special case of GMM—recovers a positive-weight average of treatment effects. In finance, hansen1997assessing show that in a misspecified asset pricing model, the GMM objective with a specific weighting matrix measures the squared error in approximating an admissible stochastic discount factor, and so measures the “size” of pricing errors in a precise sense. $\blacksquare$
@@
As our discussion of parametric models suggests, even beyond settings covered by existing theoretical guarantees, researchers’ choice of weighting matrices for GMM and related methods (e.g. simulated method of moments, minimum distance) are often informed by concerns about misspecification. To systematically explore this, we reviewed empirical papers published in the American Economic Review between 2020 and 2024 which, according to a keyword search, used moment-based methods. [^16] We excluded methods like ordinary least squares, two-stage least squares, and maximum likelihood which may be cast as special cases of GMM but which the authors did not describe as such. The search turned up 36 papers. We refer to these papers collectively as our “GMM sample.” In this sample, 22 of the papers were over-identified, and in 14 of these no results were reported based on an efficient weighting matrix. [^17] In addition, some of the papers which notionally use efficient weights plug in parameters estimated in a first step. This changes the form of the “correct” optimal weighting matrix, but is not accounted for in the implemented weighting matrices. Consequently, the number of papers in fact using theoretically efficient weights is still lower. For those models where results based on efficient GMM are not reported, the most common (non-exclusive) choices were (i) the inverse of the diagonal of the asymptotic variance-covariance matrices (4 papers) (ii) identity weighting matrix (6 papers), and (iii) user-selected weighting matrices, often explicitly chosen based on domain knowledge (7 papers).
@@
One reason the “optimal” weighting matrix may not be used in some cases is that in settings where the sample size is small relative to the number of moments, it is known that estimation of the optimal weighting matrix can lead to bias in both the estimate and the standard errors (see e.g. altonji1996small; windmeijer2005finite). One suggestion in such cases is to use the inverse of the diagonal alone, which altonji1996small found yielded “a substantial bias reduction” in their problem. [^18] Consistent with this interpretation, abebe2021selectionoftalent write that “In line with the recent literature, we use this simple weighting matrix instead of the theoretically optimal weighting matrix, which may suffer from small sample bias (altonji1996small).” However, together with optimal weighting this still covers only about half of the over-identified papers in our GMM sample.
@@
For the remaining two choices, identity weights and hand-selected weights, researchers are directly choosing the weight assigned to different moments, and thus the estimand. This is most obvious for hand-selected weighting matrices, where in many cases the authors explicitly motivate their choice of weighting matrix based on their goals for estimation. [^19] For instance, antras2020geography write that “In order to guarantee that our model provides a proper quantitative evaluation of the general-equilibrium workings of the world economy, we place a higher weight on matching the empirical moments of larger trade flows.” Similarly, einav2020screening write that “Because identification is primarily driven by the discontinuous change in screening rates at age 40, we weight more heavily moments that are closer to age 40 than moments that are associated with younger and older ages.” Perhaps more subtly, for identity weighting matrices the scale that the authors choose for each moment (e.g. whether to measure a monetary outcome in dollars or thousands of dollars) determines the weight. [^20] Weighting by the inverse of the diagonal of the variance eliminates dependence on the scale of individual moments, but is still sensitive to more general linear transformations of the moments, for instance replacing control and treatment means by control means and treatment-control contrasts. In lise2020multi_skills, for instance, the authors present a standard error formula based on identity weighting (see their Footnote 31), while their code uses a hand-selected diagonal but non-identity weighting matrix (which is consistent with the discussion in their footnote under an appropriate choice of units for the moments).
@@
A growing literature characterizes estimands obtained from misspecified models in additional settings. An early positive result by angirst1998veterans shows that in linear models where a binary treatment is as good as randomly assigned given a discrete covariate, regressions on the treatment and saturated dummies for the covariate recover a positive weighted average of causal effects. More recently, however, a large literature warns that estimands from linear (regression or instrumental variables) models fail to recover positively weighted averages of causal effects in important cases. Examples include difference in difference and event-study models estimated via two-way fixed effect regressions (chaisemartin2020twfe; callaway2020twfe; sun2021twfe), IV models estimated by two stage least squares regressions with multiple endogenous variables or insufficiently rich controls (sloczynski2020should; bhuller2024tsls; blandhol2025tsls), nonlinear estimators for linear IV, such as LIML and continuously updating GMM (kolesar2013iv; andrews2019structure), and linear regressions including both controls and multiple treatments (goldsmithpinkham2024contamination). chen2025potentialweights develops a general method for characterizing the causal estimand, if any, recovered by linear regression estimates.
Other recent papers characterize estimands in nonlinear models, including Andrews2025structural who study a class of nonlinear structural models and provide necessary and sufficient conditions for instrumental variables estimates to recover linear transformations of causal effects, as well as rambachan2025commontimeseriesestimands and kolesar2025dynamiccausaleffectsnonlinear who consider estimation of causal effects in time-series models and derive positive results for certain linear estimators, along with negative results for some popular nonlinear estimators. Other recent work extends the concept of a pseudo-true value to set-identified models (kwon2024misspecified; kaido2024informationbasedinferencemodels). $\blacksquare$